OpenStudy (anonymous):

what is the equation of a line that has a slope of 0 and goes through (5,-8)

OpenStudy (anonymous):

you can use point-slope formula which goes as y-y1=m(x+x1) the y1 being your y point, in this case would be -8. the x1 point being your x point, in this case would be 5. the m would be your slope, which is 0. now you can just plug in the numbers. Remember not to plug anything into the y and x that dont have the numbers on them y--8=0(x-5) remember two negatives make a positive in y--8. so lets do y+8=0(x-5) now, with the 0, you use the distributive property to the (x-5) 0*x = 0 and 0*5=0 now you're just left with y+8=0 subtract 8 from both sides to get the y alone, which would be y=-8